OpenStudy (anonymous):

How to simplify: 3√2x + x√8x - 5√18x I believe the first one can't be simplified further, but I'm not sure about the rest.

jimthompson5910 (jim_thompson5910):

\[\Large x\sqrt{8x}\] \[\Large x\sqrt{4*2x}\] \[\Large x\sqrt{4}*\sqrt{2x}\] \[\Large x*2*\sqrt{2x}\] \[\Large 2x\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

so \[\Large x\sqrt{8x}\] simplifies to \[\Large 2x\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

do the same for \[\Large -5\sqrt{18x}\]

OpenStudy (anonymous):

√18x = √9 * 2x = √9 * √2x = -5 * 3 √2x = -15√2x -5 -15√2x -20√2x ?

jimthompson5910 (jim_thompson5910):

good, \[\Large -5\sqrt{18x}\] simplifies to \[\Large -15\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

so \[\Large 3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}\] turns into \[\Large 3\sqrt{2x}+2x\sqrt{2x}-15\sqrt{2x}\] Now combine like terms to get \[\Large -12\sqrt{2x}+2x\sqrt{2x}\] and optionally you can factor out \(\large \sqrt{2x}\) to get \[\Large (-12+2x)\sqrt{2x}\]

OpenStudy (anonymous):

Thanks so much for all of your help!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Oh wait, so it isn't -20 it's just -15? Do you not count the 5 in the beginning?

jimthompson5910 (jim_thompson5910):

what 5 in the beginning?

jimthompson5910 (jim_thompson5910):

the 5 on \[\Large -5\sqrt{18x}\] ?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

yeah I accounted for that

jimthompson5910 (jim_thompson5910):

\[\Large -5\sqrt{18x}\] \[\Large -5\sqrt{9*2x}\] \[\Large -5\sqrt{9}*\sqrt{2x}\] \[\Large -5*3*\sqrt{2x}\] \[\Large -15\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

\[\Large -5\sqrt{18x}\] simplifies to \[\Large -15\sqrt{2x}\]

OpenStudy (anonymous):

Oh okay, I get what you're saying. Thank you! :)

jimthompson5910 (jim_thompson5910):

sure thing